An economy has a fixed price​ level, no​ imports, and no income taxes. MPC is

0.50.5​,

and real GDP is

​$250250

billion. Businesses increase investment by

​$55

billion.

Calculate the multiplier and the change in real GDP.

The multiplier is

nothing.

The increase in real GDP is

​$nothing

billion.
Answer #1

Answer

Formula :

Multiplier = 1/(1 - MPC(1 - t) + m)

where t = tax rate = 0(No taxes), m = Marginal propensity to imports = 0(as imports = 0) and MPC = Marginal propensity to consume = 0.5

Hence Multiplier = 1/(1 - 0.5(1 - 0) + 0) = 2

Hence, Multiplier = 2

Interpretation of Multiplier( = 2):

$1 change in autonomous expenditure will result in change in Real GDP by $2(multiplier).

Here Autonomous expenditure increases by $55 billion i.e. change in autonomous expenditure = 55 billion(positive means that expenditure increases). Thus real GDP will increase by 55 billion*2 = 110 billion.

Hence, the increase in Real GDP = $110 billion.